# Efficient State Preparation for Metrology and Quantum Error Correction with Global Control

### v1

@@ Abstract
In Fig. (3) we plot the infidelity of the state $\lvert\psi_{\text{f}}\rangle$ after the optimized gate protocol Eq. (3) with target states $\lvert\psi_{\text{t}}\rangle=\lvert R_{1}\rangle$ (blue with dots) and $\lvert\psi_{\text{t}}\rangle=\lvert G_{0}\rangle$ (orange with stars). We find solutions that prepare the codewords with infidelities $1-\mathcal{F}<10^{-4}$ in only $P=4$ and $P=7$ gate steps respectively. Therefore, our protocol provides a pathway towards the implementation of large $N$-qubit QEC codes utilizing only global addressability of the qubits constituting the codeword.
Considering system sizes of $N\sim$ 100, we have observed that it is possible to numerically find gate sequences with $P=\mathcal{O}(N)$ pulses with infidelities $\mathcal{O}(10^{-6})$. This suggests that there is an efficient protocol of at most a linear depth to prepare Dicke states with small infidelity.
@@
Figure 3: Infidelity to prepare the $N=9$ Ruskai $\lvert R_{1}\rangle$ and $N=13$ Gross $\lvert G_{0}\rangle$ codewords using the optimized gate sequence as a function of gate steps $P$. Despite being highly specific superpositions of Dicke states, we numerically find solutions that can prepare both target states with infidelities $1-\mathcal{F}<10^{-4}$ in $P=4$ and $P=7$ gate steps for $\lvert R_{1}\rangle$ and $\lvert G_{0}\rangle$ respectively.
@@
Realization in trapped ions: $H_{\text{OAT}}$ can be realized in trapped ions by off-resonantly driving the centre of mass (COM) mode [47, 48, 49, 38, 50]. Considering $N$ ions in a linear Paul trap, two global Raman beams detuned from the red and blue sidebands by $\delta$ leads to the following Hamiltonian in the interaction picture,
$H_{I}=\frac{\Omega\eta}{2}(ae^{-i\delta t}+a^{\dagger}e^{i\delta t})J_{z},$ (6)
@@
where we have assumed the Lamb-Dicke regime, and where $a$ ($a^{\dagger}$) are the annihilation (creation) operators of the COM mode, $\eta$ the Lamb-Dicke parameter and $\Omega$ the Rabi frequency. Note that although we require the Lamb-Dicke regime, our protocol is otherwise insensitive to temperature and therefore does not require ground state cooling [47, 51]. Evolving for time $t_{f}=2\pi/\delta$ results in a closed loop in phase space such that the motion decouples from the spin degree of freedom. The effective evolution is described by one-axis twisting, $U_{\text{OAT}}=\exp(-i\phi J_{z}^{2})$ with $\phi=\pi\eta^{2}\Omega^{2}/\delta^{2}$ the area enclosed [52]. Our optimised solutions are typically $\phi\sim 1$. Setting $\eta\Omega=2\pi\times 20$ kHz and $\delta=2\pi\times 20$ kHz results in $t_{\text{f}}=50$ $\mu$s and $\phi=\pi$. The total protocol duration $T$ is therefore $T=Pt_{\text{f}}$, and thus for all the gate sequences we present $T\lesssim 350$ $\mu$s.
@@
Off-resonant driving of the other modes will lead to non-uniform spin-dependent phases. Assuming a typical trap frequency $\omega=2\pi\times 2$ MHz, the stretch mode, which is the leading order correction, is detuned by $(\sqrt{3}-1)\omega+\delta$. The area enclosed by the stretch mode is $\sim 10^{-3}$ times smaller than the area enclosed by the COM mode, and thus for realistic parameters unwanted off-resonant driving of the collective modes is negligible compared to the effects of dephasing, which we consider below.
@@
We use a realistic dephasing rate $\gamma=2\pi\times 5$ Hz. The off-diagonal coherences of the GHZ state decay as $\sim e^{-\gamma Nt}$, such that the fidelity can be roughly estimated as $F\sim 1/2(1+e^{-\gamma Nt})$. We assume that the rotations are noise-free, as they can be performed fast relative to the spin squeezing. For the $N=9$ Ruskai and $N=13$ Gross protocols, using the $P=4$ and $P=7$ gate sequences respectively, assuming the ideal protocol prepares the state with perfect fidelity, we estimate noisy fidelities $\mathcal{F}\sim 97\%$ and $\mathcal{F}\sim 93\%$ respectively. Reducing the dephasing rate to $\gamma=2\pi\times 1$Hz increases the Ruskai and Gross codeword fidelities to $\mathcal{F}\sim 99\%$ and $\mathcal{F}\sim 98\%$ respectively. Furthermore, parametric amplification can increase the squeezing amplitude achieved in a given $t_{\text{f}}$ [53], whilst multiple Raman beatnotes can reduce off-resonant driving enabling a larger $\delta$ [54, 55]. Therefore such techniques, when combined with achievable reductions in dephasing noise, could bring state preparation infidelities of the Gross and Ruskai codewords below QEC thresholds in the near-term.
@@
Outlook: We proposed a few-gate protocol to prepare specific superpositions of Dicke states which can be easily realized across a variety of experimental platforms. Its implementation could lead to the preparation of interesting and useful entangled states, with applications in quantum metrology and quantum error correction.
@@
Finally, we note that the set of realizable interactions depends on the experimental platform, e.g. in cavity QED experiments dispersive interactions can be used to engineer the generator $\sin(J_{z}+\varphi)$ [29], whilst addressing the second red and blue sidebands in trapped ions can produce $\cosh(rJ_{z})+\cos(\varphi)\sinh(rJ_{z})$. These Hamiltonians can replace $H_{\text{OAT}}$; or the inclusion of more than one flavour of phase gate in the protocol could provide additional geometric control, potentially reducing the total number of gate steps required.
@@
Author’s note: While completing our manuscript we became aware of a recent preprint proposing a similar protocol [56].

### v2

@@ Abstract
In Fig. (4) we plot the infidelity of the state $\lvert\psi_{\text{f}}\rangle$ prepared using the optimized gate protocol Eq. (2) with target states $\lvert\psi_{\text{t}}\rangle=\lvert R_{1}\rangle$ (blue with dots) and $\lvert\psi_{\text{t}}\rangle=\lvert G_{0}\rangle$ (orange with stars). We find solutions that prepare the codewords with infidelities $1-\mathcal{F}<10^{-4}$ in only $P=4$ and $P=7$ gate steps, respectively. Therefore, our protocol provides a pathway towards the implementation of large $N$-qubit QEC codes utilizing only global addressability of the qubits constituting the codeword.
@@
Figure 4: Infidelity to prepare the $N=9$ Ruskai $\lvert R_{1}\rangle$ and $N=13$ Gross $\lvert G_{0}\rangle$ codewords using the optimized gate sequence as a function of $P$. Despite being highly specific superpositions of Dicke states, we numerically find solutions that can prepare both target states with infidelities $1-\mathcal{F}<10^{-4}$ in $P=4$ and $P=7$ gate steps for $\lvert R_{1}\rangle$ and $\lvert G_{0}\rangle$, respectively.
@@
Realization in trapped ions: $U_{\text{OAT}}$ can be realized in trapped ions by off-resonantly driving the centre of mass (COM) mode [60, 61, 62, 34, 63]. Considering $N$ ions in a linear Paul trap, two global Raman beams detuned from the red and blue sidebands by $\delta$ leads to the following Hamiltonian in the interaction picture,
$H_{I}=\Omega\eta(ae^{-i\delta t}+a^{\dagger}e^{i\delta t})J_{z},$ (5)
@@
where we have assumed the Lamb-Dicke regime, and where $a$ ($a^{\dagger}$) are the annihilation (creation) operators of the COM mode, $\eta$ the Lamb-Dicke parameter and $\Omega$ the Rabi frequency. Note that although we require the Lamb-Dicke regime, our protocol is otherwise insensitive to temperature and therefore does not require ground state cooling [60, 64]. Evolving for time $t_{f}=2\pi/\delta$ results in a closed loop in phase space such that the motion decouples from the spin degree of freedom. The effective evolution is described by one-axis twisting, $U_{\text{OAT}}=\exp(i\phi J_{z}^{2})$ with $\phi=2\pi\eta^{2}\Omega^{2}/\delta^{2}$ the area enclosed [65]. Our optimised solutions are typically $\phi\sim 1$. Setting $\eta\Omega=2\pi\times 20$ kHz and $\delta=2\pi\times 20$ kHz results in $t_{\text{f}}=50$ $\mu$s and $\phi=2\pi$. The total protocol duration $T$ is therefore $T=Pt_{\text{f}}$, and thus for all the gate sequences we present $T\lesssim 350$ $\mu$s.
@@
Here we study the effects of off-resonant driving and dephasing, with laser intensity fluctuations and non-global beams further analysed in the Supplemental Material [46]. Off-resonant driving of the other modes leads to non-uniform spin-dependent phases. Assuming a typical trap frequency $\omega=2\pi\times 2$ MHz, the stretch mode, which is the leading order correction, is detuned by $(\sqrt{3}-1)\omega+\delta$. The area enclosed by the stretch mode is $\sim 10^{-2}$ times smaller than the area enclosed by the COM mode, and thus for realistic parameters unwanted off-resonant driving of the collective modes is negligible compared to the effects of dephasing, which we consider below.
@@
We use a realistic dephasing rate $\gamma=2\pi\times 5$ Hz. The off-diagonal coherences of the GHZ state decay as $\sim e^{-\gamma Nt}$, such that the fidelity can be roughly estimated as $F\sim 1/2(1+e^{-\gamma Nt})$. We assume that the rotations are noise-free, as they can be performed fast relative to the spin squeezing. For the $N=9$ Ruskai and $N=13$ Gross protocols, using the $P=4$ and $P=7$ gate sequences respectively, assuming the ideal protocol prepares the state with perfect fidelity, we estimate noisy fidelities $\mathcal{F}\sim 97\%$ and $\mathcal{F}\sim 93\%$ respectively. Reducing the dephasing rate to $\gamma=2\pi\times 1$Hz increases the Ruskai and Gross codeword fidelities to $\mathcal{F}\sim 99\%$ and $\mathcal{F}\sim 98\%$ respectively. Furthermore, parametric amplification can increase the squeezing amplitude achieved in a given $t_{\text{f}}$ [66], whilst multiple Raman beatnotes can reduce off-resonant driving enabling a larger $\delta$ [67, 68]. Such techniques, when combined with achievable reductions in dephasing noise, could bring state preparation infidelities of the Gross and Ruskai codewords below QEC thresholds in the near-term.
@@
Outlook: We showed that global variational circuits of linear depth can prepare arbitrary superpositions of Dicke states. The circuit can be easily realized across a variety of experimental platforms. Its implementation could lead to the preparation of interesting and useful entangled states, with applications in quantum metrology and quantum error correction.
@@
Finally, we note that the set of realizable interactions depends on the experimental platform, e.g. in cavity QED experiments dispersive interactions can be used to engineer the generator $\sin(J_{z}+\varphi)$ [25], whilst addressing the second red and blue sidebands in trapped ions can produce $\cosh(rJ_{z})+\cos(\varphi)\sinh(rJ_{z})$. These Hamiltonians can replace $U_{\text{OAT}}$; or the inclusion of more than one flavour of phase gate in the protocol could provide additional geometric control, potentially reducing the total number of gate steps required.
@@
Supplemental Material: Global Variational Quantum Circuits for Arbitrary Symmetric State Preparation
I Errors due to laser intensity fluctuations
In this section we study the effect of laser intensity fluctuations on the protocol given in Eq. (2) of the main text. We focus on an implementation in trapped ions, and consider worst-case shot-to-shot noise. We assume that the noise is proportional to the one-axis twisting duration, and simulate the noise by multiplying each one-axis twisting parameter $\phi_{k}$ by a fluctuation randomly sampled from a Gaussian distribution with standard deviation $\delta\phi\times\phi_{k}$, where $\delta\phi$ is a scaling factor [69]. We obtain an average infidelity by averaging the noisy protocol infidelity over 200 realizations of the random Gaussian sampling. In Fig. S1 we plot the resulting averaged infidelity as a function of $\delta\phi$ for (a) one Haar random state for various values of $N$, and (b) the $N=9$ Ruskai and $N=13$ Gross codewords. For the sake of demonstration, the Haar random state is chosen from the $N_{\rm Haar}=200$ random states that were generated for Fig. 2 of the main text by choosing the state whose protocol parameters had the smallest total one-axis twisting amplitude, i.e. the smallest $\Phi=\sum_{i}|\phi_{i}|$. As expected, as both $N$ and $\delta\phi$ increase the infidelity increases until saturation $11-{\cal F}=$ 1 at large $\delta\phi$. We note that the infidelity of both codewords remains low, $1-\mathcal{F}\lesssim 2\times 10^{-4}$ and $1-\mathcal{F}\lesssim 2\times 10^{-3}$ for Ruskai and Gross respectively, even up to fluctuations as large as $\delta\phi=0.1\%$.
Figure S1: Averaged infidelity including shot-to-shot laser intensity fluctuations during the preparation of (a) one Haar random state and (b) Ruskai and Gross codewords. The scaling factor $\delta\phi$ enters the standard deviation $\delta\phi\times\phi_{k}$ of the Gaussian from which the noisy one-axis twisting parameter $\phi_{k}$ is sampled. The infidelity is proportional to $\delta\phi$ and $N$.
II Errors due to non-global Mølmer-Sørensen gate
In this section we consider the effect of non-global one-axis twisting. As discussed in Eq. (5) of the main text, in trapped ions one-axis twisting is realized with a Mølmer-Sørensen interaction. If the laser intensity of the Raman beam is nonuniform, the Rabi frequency $\Omega$ becomes site-dependent, so that the interaction Hamiltonian Eq. (5) becomes
$H_{I}^{\prime}(\Omega_{i})=\eta(ae^{-i\delta t}+a^{\dagger}e^{i\delta t})\sum_{i}\frac{\Omega_{i}\sigma_{i}^{z}}{2}.$ (S1)
We simulate the effect of a non-uniform Raman beam by drawing $\Omega_{i}$ from a Gaussian function of standard deviation $\sigma$ at $x$-values corresponding to the dimensionless positions of the ions $u_{i}=x_{i}/\ell$, as shown in Fig. S2(a)-(b), where the length scale is $\ell^{3}=Z^{2}e^{2}/(4\pi\epsilon_{0}M\omega^{2})$, with $Z$ the degree of ionization, $e$ the electron charge, $\epsilon_{0}$ the free space permittivity, $M$ the ion mass and $\omega$ the trap frequency. We fix $\eta=1$, $\Omega=2\pi\times 20~{}\text{kHZ}$ and vary the detuning $\delta$ so as to realize the desired one-axis twisting angle $\phi$ according to $\phi=2\pi\eta^{2}\Omega^{2}/\delta^{2}$. In panel (c) we plot the resulting infidelity as a function of $\sigma$ for the preparation of the middle Dicke state $\lvert J=N/2,0\rangle$ for $N=6$, $N=8$ and $N=10$. As expected the infidelity increases with increasing $N$, as seen in the order of magnitude increase in infidelity when going from $N=6$ to $N=10$ for a given standard deviation value $\sigma$. The infidelity also increases with the total one-axis twisting duration $\Phi=\sum_{i}\phi_{i}$, as seen in the similarities between the $N=8$ and $N=10$ curve, which have $\Phi=17.22$ and $\Phi=12.71$, respectively. Thus, finding optimal parameters that minimize the total one-axis twisting duration is crucial for minimizing the effect of non-global one-axis twisting.
Figure S2: (a) Gaussian distribution of standard deviation $\sigma$ for $\Omega_{i}$, with the dots showing the value for $\Omega_{i}$ sampled at the dimensionless positions of the ions $u_{i}$ for $N=10$. (b) Equilibrium positions for a $N=10$ ion chain. (c) Infidelity vs Gaussian standard deviation $\sigma$. The infidelity increases with increasing $N$ and increasing total one-axis twisting amplitude $\Phi$.
III Hybrid quantum-classical algorithm
In this section we provide details on how the variational circuit of Eq. (2) of the main text can be implemented as a hybrid quantum-classical algorithm [41, 44, 70, 71, 41, 42]. Specifically we consider how the cost function (infidelity) can be efficiently measured on a quantum device and then fed to the classical optimizer. The advantage of this approach is that the optimization is noise-aware, potentially enabling the mitigation of certain types of noise. For example, a naive strategy to minimize the effect of dephasing on the fidelity is to minimize the total one-axis twisting duration. The hybrid quantum-classical approach natively incorporates the effect of dephasing into the cost function by evaluating the infidelity on-device, such that the classical optimization algorithm can balance finding the (noiseless) ideal parameters with minimizing the total one-axis twisting gate time.
For the following, we assume that although the experimenter only has global control over the qubits, they have individual qubit readout, as is often the case in platforms such as trapped ions. Then, permutationally invariant (PI) tomography can be used to compute the fidelity of the prepared state with respect to the target state using $\mathcal{O}(N^{2})$ global measurement settings, where a global measurement setting is defined as a choice of single-qubit observable $A_{j}=a_{j,x}\sigma_{x}+a_{j,y}\sigma_{y}+a_{j,z}\sigma_{z}$ measured simultaneously (globally) on all $N$ qubits, i.e. $A_{j}^{\otimes N}$. Below we briefly review PI quantum tomography following Ref. [72].
Firstly, the PI density matrix is defined as
$\rho_{\rm PI}=\frac{1}{N!}\sum_{k}\Pi_{k}\rho\Pi_{k},$ (S2)
where $\Pi_{k}$ is the qubit permutation operator and $\rho$ the density matrix. A symmetric state $\rho$ is invariant under the action of $\Pi_{k}$ and thus $\rho=\rho_{\rm PI}$. To obtain $\rho_{\rm PI}$ from experimentally accessible measurements, note that $\rho_{\rm PI}$ can be written as a linear combination of $\mathcal{D}_{N}=\begin{pmatrix}N+2\\ N\end{pmatrix}=\frac{1}{2}(N^{2}+3N+2)$ operators of the form $(\sigma_{x}^{\otimes k}\sigma_{y}^{\otimes l}\sigma_{z}^{\otimes m}\mathbb{1}^{\otimes n})_{\rm PI}$, where $k+l+m+n=N$ and where the $(\cdot)_{\rm PI}$ notation indicates an operator of the form of Eq. (S2). To avoid measuring $(\sigma_{x}^{\otimes k}\sigma_{y}^{\otimes l}\sigma_{z}^{\otimes m}\mathbb{1}^{\otimes n})_{\rm PI}$ directly which would require single-qubit control, note the decomposition [72]
$\langle(\sigma_{x}^{\otimes k}\sigma_{y}^{\otimes l}\sigma_{z}^{\otimes m}\mathbb{1}^{\otimes n})_{\rm PI}\rangle=\sum_{j}^{\mathcal{D}_{N}}c_{j}^{(k,l,m)}\langle(A_{j}^{\otimes(N-n)}\otimes\mathbb{1}^{\otimes n})_{\rm PI}\rangle,$ (S3)
where $c_{j}^{(k,l,m)}$ are real coefficients and $A_{j}^{\otimes(N-n)}$ the $j$-th measurement setting acting on $(N-n)$ qubits. The choice of coefficients $c_{j}^{(k,l,m)}$ and measurement settings $A_{j}$ is not arbitrary. The optimal choices can be found using numerical optimization techniques as detailed in Ref. [72].
Next, observe that the choice of measurement setting $A_{j}$ appearing in the RHS of Eq. S3 is independent of $k$, $l$ and $m$. For a given $j$, all of the expectation values $\langle(A_{j}^{\otimes(N-n)}\otimes\mathbb{1}^{\otimes n})_{\rm PI}\rangle$ for $0\leq n\leq N$ can be reconstructed in post-processing from the coincidence counts of the global measurement setting $A_{j}^{\otimes N}$. The quantities to be measured appearing on the right-hand side are therefore entirely independent of the coefficients $k$, $l$, $m$ and $n$ appearing on the left-hand side.
As such, for all values of $k,l,m$ and $n$, $\langle(\sigma_{x}^{\otimes k}\sigma_{y}^{\otimes l}\sigma_{z}^{\otimes m}\mathbb{1}^{\otimes n})_{\rm PI}\rangle$ can be reconstructed from the coincidence counts of an optimized set of $\mathcal{D}_{N}$ global measurement settings $\{A_{j}^{\otimes N}\}_{j=1}^{\mathcal{D}_{N}}$ with optimized coefficients $c_{j}^{(k,l,m)}$. Therefore, $\rho_{\rm PI}$ can be obtained using $\mathcal{D}_{N}$ global measurement settings.
Finally, we note that in the case of leakage from the PI density matrix, i.e. due to any non-global operations or decoherence sources, the fidelity between the target PI state $\rho_{\rm PI}$ and the experimentally prepared state $\rho$ can be efficiently lower bounded [72].
III.1 Fidelity proxy
Although the number of measurement settings required to perform PI tomography scales efficiently (quadratically) with $N$, here we discuss how under certain assumptions, a proxy for the fidelity cost function can be obtained using only a single measurement setting. Firstly, we assume that the target state $\lvert\psi_{t}\rangle$ can be written in the $J=N/2$ Dicke state basis with real coefficients,
$\lvert\psi_{t}\rangle=\sum_{i=-J}^{J}b_{i}\lvert J,i\rangle,\qquad b_{i}\in\mathbb{R}.$ (S4)
Many states of interest are of the form of Eq. (S4), including any single Dicke state and both Gross and Ruskai codewords [7, 6]. For a given Dicke state $\lvert J,m\rangle$ we denote its corresponding density matrix $\rho_{J,m}=\lvert J,m\rangle\langle J,m\rvert$, which has a one at position $(m+N/2+1,m+N/2+1)$ and zeros elsewhere.
Next, we define a proxy for the fidelity,
$\mathbb{F}(\rho_{\rm exp},\rho_{t})=\sum_{i=-J}^{J}\operatorname{Tr}(\rho_{\rm exp}\rho_{J,i})\operatorname{Tr}(\rho_{t}\rho_{J,i}),$ (S5)
where $\rho_{\rm exp}$ is the experimentally prepared state and $\rho_{t}=\lvert\psi_{t}\rangle\langle\psi_{t}\rvert$ the target state. In words, $\mathbb{F}$ measures the inner product between the vectors $(\operatorname{Tr}(\rho_{\rm exp}\rho_{J,i=-J}),\ldots,\operatorname{Tr}(\rho_{\rm exp}\rho_{J,i=J}))$ and $(\operatorname{Tr}(\rho_{t}\rho_{J,i=-J}),\ldots,\operatorname{Tr}(\rho_{t}\rho_{J,i=J}))$ which correspond to the decomposition of $\rho_{\rm exp}$ and $\rho_{t}$ on the Dicke state basis. Although $\mathbb{F}$ does not provide any information about the overall coherence of the quantum state, it does give information about our ability to coherently prepare a specific Dicke state $\lvert J,i\rangle$ (which in the computational basis can be highly entangled) with target population $|b_{i}|^{2}$. Using the fact that $\rho_{J,m}$ can be written as a sum over powers of $J_{z}$ as $\rho_{J,m}=\sum_{i=0}^{N}c_{i,m}J_{z}^{i}$ for coefficients $c_{i,m}\in\mathbb{R}$ (for proof see below), the fidelities of $\rho_{\rm exp}$ with respect to each $\rho_{J,m}$ can be written as
$\operatorname{Tr}(\rho_{\rm exp}\rho_{J,m})=\sum_{i=0}^{N}c_{i,m}\operatorname{Tr}(J_{z}^{i}\rho_{\rm exp}).$ (S6)
The expectation values $\operatorname{Tr}(J_{z}^{i}\rho_{\rm exp})$ can in turn be decomposed using [73]
$J_{\alpha}^{n}=\sum_{i=0}^{n}d_{i,\alpha}\sum_{k}\left[\Pi_{k}\sigma_{\alpha}^{\otimes i}\otimes\mathbb{1}^{\otimes(N-i)}\Pi_{k}\right],\qquad\alpha\in\{x,y,z\},$ (S7)
where the summation over $k$ is over all permutations of the qubit ordering, and where the coefficients $d_{i,\alpha}$ are defined by the expansion. For example, for $\alpha=x,n=3$ we have $d_{0,1,2,3,\alpha}=0,7/8,0,6/8$ because $J_{x}^{3}=\frac{1}{8}(\sigma_{x}^{(1)}+\sigma_{x}^{(2)}+\sigma_{x}^{(3)})^{3}=\frac{6}{8}\sigma_{x}^{(1)}\sigma_{x}^{(2)}\sigma_{x}^{(3)}+\frac{7}{8}(\sigma_{x}^{(1)}+\sigma_{x}^{(2)}+\sigma_{x}^{(3)})$ where $\sigma_{x}^{(i)}$ acts on the $i$th qubit.
The RHS of Eq. (S7) can always be reconstructed from the coincidence counts of a single measurement setting $\sigma_{z}^{\otimes N}$. Thus the populations $\operatorname{Tr}(\rho_{\rm exp}\rho_{J,m})$ for all $m\in\{-J,\dots,J\}$, and therefore fidelity proxy $\mathbb{F}$ of Eq. (S5) can be obtained from a single measurement setting $\sigma_{z}^{\otimes N}$.
A limitation of $\mathbb{F}$ is that, although it provides information about our ability to coherently prepare the state $\lvert J,i\rangle$ with weight $|b_{i}|^{2}$, it does not give any information about the coherence of the overall superposition. For example, perhaps instead of preparing the coherent superposition $\lvert\psi\rangle=1/\sqrt{|b_{k}|^{2}+|b_{l}|^{2}}(b_{k}\lvert J,k\rangle+b_{l}\lvert J,l\rangle)$, we prepared the mixed state $\rho=|b_{k}|^{2}\lvert J,k\rangle\langle J,k\rvert+|b_{l}|^{2}\lvert J,l\rangle\langle J,l\rvert$. One straightforward strategy to mitigate this is to use $\mathbb{F}$ during an initial optimization step before switching to full PI tomography, or to use full PI tomography as a verification of the parameters optimized using $\mathbb{F}$.
Proof that $\rho_{J,m}=\sum_{i=0}^{N}c_{i,m}J_{z}^{i}$. First, note that in the basis of the Dicke states $\{\lvert J=N/2,i=-J\ldots J\rangle\}$, $\rho_{J,m}$ is always a diagonal matrix with a one at position $(m+N/2+1,m+N/2+1)$. Because $J_{z}$ is diagonal, ${\rm diag}(J_{z})=[-N/2,-N/2+1,\dots,N/2-1,N/2]$ and its powers are simply ${\rm diag}(J_{z}^{i})=[(-N/2)^{i},(-N/2+1)^{i},\dots,(N/2-1)^{i},(N/2)^{i}]$, with every $J_{z}^{i}$ for $0\leq i\leq N$ linearly independent. The equation $\rho_{J,m}=\sum_{i=0}^{N}c_{i,m}J_{z}^{i}$ can be vectorized into a set of $N+1$ linear equations with $N+1$ variables by defining $x=\{c_{i,m}\}_{i=0}^{N}$, $b={\rm diag}(\rho_{J,m})$ and the $(N+1)\times(N+1)$ coefficient matrix $A=\{{\rm diag}(J_{z}^{i})\}_{i=0}^{N}$, such that the equation reduces to standard form $Ax=b$. Because the columns of $A$ are each from a given $J_{z}^{i}$ for $0\leq i\leq N$, all columns are linearly independent. Thus $A$ is invertible, yielding the unique solution $x=A^{-1}b$.